OpenStudy (anonymous):

Determine the number and type of solutions for 5x2 + 1x + 7 = 0.

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@sami-21

Parth (parthkohli):

Do you know what 'discriminant' is?

OpenStudy (anonymous):

nope

Parth (parthkohli):

Given that an expression is \(ax^2 + bx + c\), then the discriminant is \(b^2 - 4ac\)

hero (hero):

I just told it to you two seconds ago

Parth (parthkohli):

Do you have any idea of what \(a,b,c\) are here in this case?

OpenStudy (anonymous):

@hero....umm..i forget...remind me? :D

Parth (parthkohli):

\[\Delta = b^2 - 4ac \]

hero (hero):

lol and Parth just told it to you just now.

Parth (parthkohli):

lol

hero (hero):

\(D = b^2 - 4ac\)

OpenStudy (anonymous):

k now that i know what a discriminant is...anyone mind telling me how to solve this... :)

hero (hero):

@yummydum, keep repeating this to yourself \(D = b^2 - 4ac\)

Parth (parthkohli):

If D > 0, then we have two real solutions. If D = 0, then we have one real solution. If D < 0, then we have no real solutions.

OpenStudy (anonymous):

\[D=-139\]

OpenStudy (anonymous):

no real solutions...?

Parth (parthkohli):

Yep.

Parth (parthkohli):

So, we have two imaginary solutions.

OpenStudy (anonymous):

1 Rational Double 2 Irrational 2 Rational 2 Complex these are my choices...

Parth (parthkohli):

Complex = Imaginary.

OpenStudy (anonymous):

irrational?

OpenStudy (anonymous):

oh...okay :D

Parth (parthkohli):

Yeah. :)

OpenStudy (anonymous):

thanks :)